OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, -2) and a directrix at y = 2.

OpenStudy (anonymous):

@ganeshie8 @Loser66

OpenStudy (anonymous):

least favorite problems lol, mind if I help?

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

standard form: y = a(x-h)^2 + k

OpenStudy (anonymous):

BTW did my help earlier help you?

OpenStudy (anonymous):

yes quite a bit actually, sorry i didn't help so much I was a bit confused. but I will make it up to you now

OpenStudy (anonymous):

 lol kind of failed but you get the point. so the vertex is halfway between these points. so the vertex is actually at (0,0) in the standard from (h,k) is the vertex following so far?

OpenStudy (anonymous):

Oh no its fine I understand!

OpenStudy (anonymous):

y = a(x-h)^2 + k y = a(x-0)^2 + 0 y = a(x)^2 a = 1/4c c is 2, the difference between the focus and the vertex so a = 1/(4*2) a = 1/8 y = 1/8 (x)^2

OpenStudy (loser66):

Focus (0,-2) hence p=-2, right? So just plug it in x^2=4py

OpenStudy (anonymous):

@Loser66 hey my final answer is y = 1/8(x)^2 Is that right?

OpenStudy (loser66):

Same

OpenStudy (anonymous):

awesome so do you get that @PHUNISH

OpenStudy (anonymous):

Oh okay thanks guys1

OpenStudy (anonymous):

no problem

OpenStudy (loser66):

But, to focus, direct fix oinformat6, you should apply the formula